OpenStudy (anonymous):

An investor wants to analyze the earnings of a mutual fund account. Seven years ago, the value of the account was $17,000 and it is now worth $21,250 (no additional deposits were made). If the account is compared to a bank account paying interest that is compounded continuously, what interest rate would the bank account have to pay to match the mutual fund accounts earnings?

OpenStudy (anonymous):

A=Pe^(rt) t=7 P= 17000 A=21250 just solve for r and replace the values

OpenStudy (anonymous):

but what about "m"? isn't the formula for it A=P(1+r/m) ^m*t

OpenStudy (anonymous):

I can't quite figure out what number to put for "continuous compounding".

OpenStudy (anonymous):

the problem says compounded continuously, so the formula is A=Pe^(rt) this other formula A=P(1+r/m) ^m*t is for interest compounded

OpenStudy (anonymous):

OH!!

OpenStudy (jdoe0001):

$$ A(t)=Pe^{rt}\\ A= principal= 21250\\ P17000\\ t=7\\ 21250=17000e^{7r}\\ --------------------\\ ln(21250)= \color{blue}{ln(17000e^{7r}) \implies ln(17000)+ln(e^{7r})}\\ ln(21250)-ln(17000)=7r $$

OpenStudy (jdoe0001):

I think you can get "r" from there :|

OpenStudy (jdoe0001):

I meant to say P=17000, btw :|

OpenStudy (anonymous):

so if it said it was a daily compound you would use the formula I had put originally and "m" would be 365, right?

OpenStudy (jdoe0001):

@hvb91 then the 'day' would be the 'period', and for a 365 lifetime, yes

OpenStudy (jdoe0001):

a period can be any arbitrary term really, 2 months, a year, a month